OpenStudy (anonymous):

Please Help, Two cars leave town going in opposite directions. One car is traveling 55 miles per hour (mph), and the other car is traveling 40 mph. How long will it take before the cars are 285 miles apart?

OpenStudy (dls):



OpenStudy (dls):

R(ab)=55+40=95mph Distance=285miles Time=285/95=>3

OpenStudy (anonymous):

Thanks ! :)